In each of the cases that follow, the components of a vector A? are given. Use trigonometry to find the magnitude of that vector and the counterclockwise angle it makes with the +x axis.


Part A

Ax = 3.00m ,  Ay = 7.00m

Part B
Ax = 3.00m ,  Ay = 7.00m
Please give your answer in the interval [0,360]?.
Answer #1

magnitude = sqrt(3^2+7^2)= 7.61


theta= arctan(7/3)= 66.8 degrees


part A and B are exacly the same